A 200N block rests on an inclined plane at 30° to the horizontal. A force parallel to and acting up the plane is applied to the body if coefficient is 0.20. Find the value of force
1
Expert's answer
2019-02-04T09:54:41-0500

There are three forces that act on the block: the force of gravity (or weight)

"mg"

directed downward and can be resolved into two perpendicular components (

"F_{||} = mgsin\alpha"

and

"F_{\perp} = mgcos\alpha"

), the force of reaction

"N"

directed perpendicular to the surface, the applied force

"F_p"

directed parallel and upward to the plane. Let’s draw a free-body diagram and write all forces that act on a block:

There are three forces that act on the block: the force of gravity (or weight)

"mg"

directed downward and can be resolved into two perpendicular components (

"F_{||} = mgsin\alpha"

and

"F_{\perp} = mgcos\alpha"

), the force of reaction

"N"

directed perpendicular to the surface, the applied force

"F_p"

directed parallel and upward to the plane. Let’s draw a free-body diagram and write all forces that act on a block:

Since the block is at rest the applied force is equal to the rolling force:

"F_p - F_r = 0,""F_p - mgsin\alpha = 0,""F_p = mgsin\alpha = Wsin\alpha."

From this formula, we can find the applied force:

"F_p = Wsin\alpha = 200 N \cdot sin30^\circ = 100 N."

Answer:

"F_p = 100 N"

.